OpenStudy (anonymous):

find the indicated outputs for f(x) = 5xto second power - 2x f(0) = --? f(-1) = --? f(2) = --?

OpenStudy (anonymous):

find the indicated outputs for f(x) = 5xto second power - 2x f(0) = --? f(-1) = --? f(2) = --?

OpenStudy (anonymous):

\[f(x)=5x^2-2x\] \[f(o)=0\] \[f(-1)=5(-1)^2-2\times -1=5+2=7\]

OpenStudy (anonymous):

\[f(2)=5(2)^2-2\times 5=5\times 4-4=20-4=16\]